OpenStudy (anonymous):

Simplify: 5 sq rt 7 -3 square root 5 - 6 square root 7 a. -4 square root 19 b. -22 c. -square root 7 - 3 square root 5 d. -1-3 square root 5

OpenStudy (anonymous):

\[5\sqrt{7} - 3\sqrt{5} - 6\sqrt{7} = -\sqrt{7} - 3\sqrt{5}\]